# The Theory of Economic Complexity

### v1

@@ 5 Prices, Wages, and Consumption
dividing the numerator and denominator by $1/N_{p}$ (one over the total number of activities) we can transform the sums into averages to obtain an equilibrium wage $w_{c}^{*}$:
$w^{*}_{c}=\frac{N_{p}(\langle\pi\rangle+\langle q\pi\rangle(r_{c}-1))}{L_{c}}$ (56)
@@
This finding is consistent with the notion that economic complexity, which we now understand as an estimate of $r_{c}$, implies an equilibrium level of wages for an economy, and thus, explains future economic growth. In this model, economies must have a wage given by eqn. 56 in equilibrium. When out of equilibrium, economies should adjust (to first order) according to:
@@
That brings us again to a specialization condition based on two clusters where economies with an above average probability of being endowed with the capability ($r_{c}>\langle r\rangle$) are specialized in products with a higher probability of requiring the capability, and where those with a below average probability of being endowed with the capability (($r_{c}<\langle r\rangle$))specialize in less demanding products. Yet, the threshold for activities is now:
@@
which after some algebra can be brought to the form161616here we used the notion that averages are constants to arrive to an expression where $\pi_{p}$ is expressed as a function of its ensemble averages.:
@@ 6 Relatedness and The Product Space
The other key observable used frequently in the economic complexity literature is a network connecting related activities [47, 48, 50, 53, 51, 123, 124, 125, 126, 86, 127, 88, 89, 128, 129, 130, 131, 52, 2]. When these activities are products, this network goes by the name of "product space." From an application perspective, the product space is used to estimate the potential of economy in an activity (e.g. the probability that a city specializes in an industry [48, 49, 50, 88], a country starts exporting a product [47, 53, 2], or a university starts producing papers in a given field [51, 52]. These estimates of potential are known as measures of relatedness, and are akin to traditional recommender system methods in computer science [132]. Yet, in the economic complexity literature, they are used to explain economic development trajectories (e.g. countries entering new products) instead of individual consumption patterns (e.g. customers choosing to purchase a products at an online retailer) or to explore strategies to optimize industrial promotion efforts[133, 134, 135] [^17] In recent years there have also been multiple efforts to look at relatedness in the context of sustainability, starting from the idea of a green product space, [136, 137, 138, 139, 140, 141, 142, 143]
@@
Here we begin by focusing on a particular characteristics of the product space that was emphasized when it was introduced as a network nearly twenty years ago: the fact that the core of the product space, its most densely connected part, is composed of high-complexity activities[47].
This is a characteristic that is true for networks derived from trade data, since networks derived from other data can have different forms. For example, networks connecting research fields based on citation patterns or co-authorships tend to follow a "ring" structure [51, 67]. Networks connecting skills based on the occupations that require them tend to follow a "dumbell" structure (two big clusters connected by a bridge) [123].
Figure 14: Average correlation between $ECI$ and the average capability endowment of an economy $\langle r\rangle_{c}$ as a function of the mixing probability $\pi$ (number indicated on the top of each chart). We can see that the correlation between $ECI$ and $\langle r\rangle_{c}$ remains close to one for mixing probabilities above 0.4.
We begin our exploration by of the structure of the product space implied by the single and multi-capability theory by estimating a measure of proximity, which is an estimate of the similarity between products. Unlike in the case of $ECI$, where we have a more strict definition based on a second eigenvector, measures of proximity, in both the economic complexity and recommender systems literature, tend to be more ad-hoc, since there are many ways to estimate similarity among pairs of activities. In [47], proximity was introduced using the minimum of the conditional probability that two products are exported in tandem. In our notation, this translates to:
$\phi_{pp^{\prime}}=\frac{\sum_{c}M_{cp}M_{cp^{\prime}}}{\text{max}(Mp,M_{p^{\prime}})}$ (78)
In [48] they use simply the number of activities that are common to two economies.
@@
In general, it is not uncommon to find proximity matrices and recommender systems based on variations of $\sum_{c}M_{cp}M_{cp^{\prime}}$ (usually with a normalization), so we will begin by exploring this basic form.
@@
$\phi_{pp^{\prime}}=\sum_{c}M_{cp}M_{cp^{\prime}}=M_{p}\quad if\quad q_{p}>\langle q\rangle\quad\&\quad q_{p^{\prime}}>\langle q\rangle$ (80)
$\phi_{pp^{\prime}}=\frac{\sum_{c}M_{cp}M_{cp^{\prime}}}{\text{max}(M_{p},M_{p^{\prime}})}=1\quad if\quad q_{p}>\langle q\rangle\quad\&\quad q_{p^{\prime}}>\langle q\rangle$ (81)
@@
A more interesting exercise is to consider the networks implied by the multi-capability model. Here we present three examples in which we estimate networks for different model parameters that we visualize by estimating their minimum spanning tree and adding on top of that all of the links that are one standard deviation above the mean. This is a similar visualization exercise than the one used in the paper that introduced the product space network.
Figure 14 presents this exercise for a model involving 200 activities, 100 economies, and 10 capabilities. The color of the nodes indicates the complexity of the activity (with darker nodes being higher complexity). The number on top of each network visualization shows the mixing parameter $\pi$ used to combine random and non-random capabilities.
We can see clearly in this example that all of the networks that are above the phase transition threshold are centered around a core of high-complexity activities, with lower complexity activities being peripheral. This reproduces the empirical observation presented in the original product space paper, which claimed that the core of the product space is composed of more sophisticated activities.
@@
But can we use this model to generate the network observed for research activities, which follows a ring instead of a core-periphery structure? Or do we need to radically change our assumptions to obtain that shape?
@@
Here, we use a parametrization where we combine a symmetric Toeplitz circulant matrix and a random matrix by using proportions of ($\alpha$) and ($1-\alpha$). A circulant matrix is a particular type of Toeplitz matrix that has periodic boundary conditions. A symmetric circulant matrix can be constructed by starting from a row that is symmetric with respect to the center. Here, we generate symmetric circulant matrices using linearly spaced probabilities for $r_{c}$ and $q_{p}$ that grow symmetrically from the center column of the first row. Figure 15 shows an example of this parametrization for a model with 200 economies, 200 activities, and 200 capabilities. We note that in this model talking about higher and lower complexity economies is not a useful construct, since economies do not differ on their average capability endowment (they are all equal on average), but in which subset of capabilities they are specialized in.
Figure 16 shows the network derived form this parametrization, visualized using the same method than before (minimum spanning tree, plus links that are one standard deviation above the average weight). The visualization shows a clear ring structure mimicking the one observed in networks involving research fields. The connectivity pattern of this network can be interpreted as research fields having a few related activities that share capabilities among them (e.g. capabilities are more re-deployable between molecular biology and biochemistry, than between polymer sciences and experimental psychology). This results in a network structure where each field is connected to a few neighbors.
Finally, we use the same approach to model a “dumbbell” network, which is a network with two well-defined yet clusters, such as the one observed when connecting skills and occupations [123]. Figures 17 and 18 show an example with 100 economies, 500 activities, and 20 capabilities. We note that obtaining this dumbbell structure requires a good level of mixing between the clusters, which can be achieved by setting the noise levels to be high enough so that some of the between cluster links are comparable in strength to the withing cluster links.
What is exciting about this general idea, is that it provides us with an intuitive way to map capability endowments to network structures. For instance, the core periphery-structure of the product space suggest that the capabilities associated with exporting products are correlated among economies, with high complexity economies like that of Singapore, Japan, or the United States, having high-values across a wide set of capabilities. The ring structure of the research space tells a different story. It is a story of specialization in a world of fine grained capabilities. Similarly, we can use this intuition to think about dumbbell structures, which can be modeled by assuming capability endowments made of slightly overlapping blocks.
@@ 7 Conclusion
In this paper, we provided an analytical foundation for the economic complexity index ($ECI$) starting from a production function. For the single-capability model, we were able to derive its key eigenvectors analytically and show that $ECI$ is a statistic separating economies among those with an above- and below-average probability of having a capability. We then extended this result numerically to a multi-capability setting to show that $ECI$ is as a monotonic estimator of an economy’s average capability endowment—even when a substantial share of the capabilities are randomly assigned. In the multi-capability model, $ECI$ is no longer a discrete measure separating low from high capability economies, but a monotonic transformation of the average capability endowment of an economy. These findings differentiate $ECI$ from measures of diversity, which peak for capability endowments below the maximum (they are non-monotonic functions of $r_{c}$), and thus are non-ideal estimates of the complexity of an economy. These results validate $ECI$ as a measure of composition or complexity, since they show the eigenvector captures information about an economy being endowed with multiple capabilities, regardless of how these capabilities are defined.
@@
Interestingly, our main result does not depend on assuming an stochastic model or a theory based on capabilities, since the basic idea can be easily generalized to models including factors that are specific to economies and activities. The key condition for the measure of complexity to work is for output to not be perfectly proportional to factor endowments. This condition can be achieved by simply shifting the production function by a constant to make it non-multiplicatively separable. [^18] This mechanism is akin to the idea of symmetry-breaking in physics, since the shift removes symmetries of the function. For example $K_{c}^{\gamma}$ satisfies the scale-invariance symmetry $f(\lambda K)=\lambda^{\gamma}f(K)$, whereas $B+K_{c}^{\gamma}$ does not have this symmetry.
@@
What is also interesting is that the condition needed for our main result to hold comes from calculating the matrix of specialization $R_{cp}$. This is a key difference with previous attempts to connect economic complexity theory and empirics [1, 54] which jumped directly to the binary specialization matrix $M_{cp}$. That assumption results in a monotonic relationship between the number of activities an economy is specialized (its diversity) and its capability endowments [^19] That equation is provided in [54]., which is an uncomfortable result since the empirical work had shown that measures of diversity fail to explain future economic growth like measures of complexity do [1]. We now understand that a key step in the calculation is estimating the specialization matrix, and that skipping this step in theoretical work results in a flawed connection. This change only helps provide a tight connection between the capability model and the economic complexity index, but also explains other findings, like that of Imbs and Warcziag [81], which says that economies diversify only until a certain point. It also opens questions about alternative measures of complexity. During the last fifteen years, many alternatives to the economic complexity index have been proposed, such as the Fitness index [11], the Ability index [12], and several others[79, 80, 148, 13, 149, 150, 151, 61]. Since these indexes tend to exhibit strong correlations with $ECI$, our results provide a way to theoretically explore whether they are also monotonic functions of an economy’s capability endowment ($r_{c}$ and $\langle r_{c}\rangle$), and if they are, open the question about the importance of the functional form connecting this two quantities.
@@
Our work also embedded this model in a short-run equilibrium framework including wages, consumption, and prices. Overall, we obtain reasonable results for all of them. Wages increase with capability endowments and prices are higher for more demanding products. Interestingly, prices do not affect the specialization condition, meaning that they leave the connection between capability endowments and economic complexity unchanged. [^20] We assume prices depend only on the activities and are the same across economies.
@@
Together, these findings help resolve a few long-standing tensions in the economic complexity literature. First, and most importantly, the disconnect between its empirical metrics and their theoretical underpinnings. Our findings show that $ECI$ is not an arbitrary or ad-hoc measure, but can be thought of as an estimator of an economy’s capability endowments derived from its pattern of specialization. This is an interesting finding, since it provides a mean to estimate the combined presence of factor or capabilities even when these cannot be identified [^21] This has for long been a tenet of this literature [47, 1, 106]. For example, in [47] relatedness is introduced as an “agnostic approach [$\ldots$] based on the idea that, if two goods are related because they require similar institutions, infrastructure, physical factors, technology, or some combination thereof, they will tend to be produced in tandem,” and in [106] complexity metrics are described as “an estimate of the overall potential of an economic structure, based on information on the geographic distribution of economic activities. As such, they help anticipate economic growth or emissions, not by identifying a specific factor, but by estimating their combined presence.”
@@
And third, we provide theoretical underpinnings for the structure of the networks of related activities. Our findings show that the structure of these networks reflect how capabilities are distributed across economies and activities.
@@ Acknowledgments
This work owes a very special acknowledgment to Cristian Jara-Figueroa. In 2014, Cristian joined my (César’s) group at the MIT Media Lab. During the first year of his Master’s he worked on the mathematical theory of economic complexity producing an impressive internal manuscript with many results. Those results were never published, but they stayed with my group. Eleven years later, in 2025, while looking at Cristian’s work, I realized we had made an important and simple mistake at the very beginning, which was to assume that the capability model was a model of $M_{cp}$ instead of a model of the output matrix $Y_{c,p}$. This changed everything and motivated me to go back to square one to start estimating the intermediate matrices in the model (such as $R_{cp}$). In my mind this work owes enormously to that effort by Cristian many years ago. We would also like to acknowledge comments by Johanness Wachs and other members of the Center for Collective Learning. The section on prices and wages was motivated by a very useful conversation with Jean Tirole.

### v5

@@ 6 Prices, Wages, and Consumption
dividing the numerator and denominator by $1/N_{p}$ (one over the total number of activities) we can transform the sums into averages to obtain an expected wage $w_{c}^{*}$:
$w^{*}_{c}=\frac{N_{p}\left(\langle\pi\rangle+\langle q\pi\rangle(r_{c}-1)\right)}{L_{c}}$ (60)
@@
This finding is consistent with the notion that economic complexity, which we now understand as an estimate of $r_{c}$, implies an equilibrium level of wages for an economy, and thus, explains future economic growth. In this model, economies must have a wage given by eqn. 60 in equilibrium. When out of equilibrium, economies should adjust (to first order) according to:
@@
That brings us again to a specialization condition based on two clusters where economies with an above average probability of being endowed with the capability ($r_{c}>\langle r\rangle$) are specialized in products with a higher probability of requiring the capability, and where those with a below average probability of being endowed with the capability ($r_{c}<\langle r\rangle$) specialize in less demanding products. Yet, the threshold for activities is now:
@@
which after some algebra can be brought to the form222222here we used the notion that averages are constants to arrive to an expression where $\pi_{p}$ is expressed as a function of its ensemble averages.:
@@ 7 Relatedness and The Product Space
Another observable used frequently in the economic complexity literature is the network of related activities [1, 3, 5, 83, 6, 91, 129, 130, 131, 87, 132, 89, 90, 133, 134, 135, 136, 7, 2]. When these activities are products, this network goes by the name of "product space." From an application perspective, the product space is used to estimate the potential of economy in an activity (e.g. the probability that a city specializes in an industry [3, 4, 5, 89], a country starts exporting a product [1, 83, 2], or a university starts producing papers in a given field [6, 7]. These estimates of potential are known as measures of relatedness, and are akin to traditional recommender system methods in computer science [137]. Yet, in the economic complexity literature, they are used to explain economic development trajectories (e.g. countries entering new products) instead of individual consumption patterns (e.g. customers choosing to purchase a products at an online retailer) and to explore strategies to optimize industrial promotion efforts[138, 139, 140] [^23] In recent years there have also been multiple efforts to look at relatedness in the context of sustainability, starting from the idea of a green product space, [141, 142, 143, 144, 145, 146, 147, 148]
@@
Here we begin by focusing on a particular characteristics of the product space that was emphasized when it was introduced nearly twenty years ago: the fact that the core of the product space, its most densely connected part, is composed of high-complexity activities[1].
This is a characteristic that is particular to networks derived from trade data, since networks derived from other sources can have different forms. For example, networks connecting research fields based on citation patterns or co-authorships tend to follow a "ring" structure [6, 68]. Networks connecting skills based on the occupations that require them tend to follow a "dumbell" structure (two big clusters connected by a bridge) [91].
Figure 15: Networks of related activities generated for models with 200 activities, 100 economies, and 10 capabilities. Node shading corresponds to activity complexity (darker nodes denote greater complexity). Each panel is labeled with the value of the mixing parameter $\pi$, which specifies the degree to which random and non-random capabilities are combined (Equation (43)).
We begin our exploration of the structure of the product space by estimating a measure of proximity, which is an estimate of the similarity between products. Unlike in the case of $ECI$, where we have a more strict definition based on a second eigenvector, measures of proximity, in both the economic complexity and recommender systems literature, tend to be more free. In [1] for instance, proximity was introduced using the minimum of the conditional probability that two products are exported in tandem. In our notation, this translates to:
$\phi_{pp^{\prime}}=\frac{\sum_{c}M_{cp}M_{cp^{\prime}}}{\text{max}(M_{p},M_{p^{\prime}})}$ (82)
In [3] proximity is simply the number of activities that are common to two economies.
@@
In general, it is not uncommon to find proximity matrices and recommender systems based on variations of $\sum_{c}M_{cp}M_{cp^{\prime}}$ (usually with a normalization), so we will focus our exploration in this basic form.
@@
$\phi_{pp^{\prime}}=\sum_{c}M_{cp}M_{cp^{\prime}}=M_{p}\quad\text{if}\quad q_{p}>\langle q\rangle\quad\&\quad q_{p^{\prime}}>\langle q\rangle$ (84)
$\phi_{pp^{\prime}}=\frac{\sum_{c}M_{cp}M_{cp^{\prime}}}{\text{max}(M_{p},M_{p^{\prime}})}=1\quad\text{if}\quad q_{p}>\langle q\rangle\quad\&\quad q_{p^{\prime}}>\langle q\rangle$ (85)
@@
A more interesting exercise is to consider the networks implied by the heterogeneous multi-capability model. Here we present three examples in which we estimate networks for different model parameters that we visualize by estimating their minimum spanning tree and adding on top of that all of the links that are one standard deviation above the mean. This is a similar visualization exercise than the one used in the paper that introduced the product space network and will ensure the same visualization method is used for all examples.
Figure 15 presents this exercise for a model involving 200 activities, 100 economies, and 10 capabilities. Darker nodes indicate higher complexity. The number on top of each network visualization shows the mixing parameter $\alpha$ used to combine random and non-random capabilities.
We can see clearly in this example that all of the networks that are above the phase transition threshold are centered around a core of high-complexity activities, with lower complexity activities being peripheral. This reproduces the empirical observation that the core of the product space is composed of more sophisticated activities.
@@
But can we use this model to generate the network structure of the research space, which follows a ring instead of a core-periphery structure? Or do we need to radically change our assumptions to obtain that shape?
@@
Here, we use a parametrization where we combine a symmetric Toeplitz circulant matrix and a random matrix by using proportions of ($\alpha$) and ($1-\alpha$). A circulant matrix is a particular type of Toeplitz matrix that has periodic boundary conditions. A symmetric circulant matrix can be constructed by starting from a row that is symmetric with respect to the center. Here, we generate these circulant matrices using linearly spaced probabilities for $r_{c}$ and $q_{p}$ that grow symmetrically from the center column of the first row. Figure 16 shows an example of this parametrization for a model with 200 economies, 200 activities, and 200 capabilities. We note that in this model talking about higher and lower complexity economies is not a useful construct, since economies do not differ on their average capability endowment (they are all equal on average), but in which subset of capabilities they are specialized in.
Figure 17 visualizes the network derived form this parametrization using the same method than before (minimum spanning tree, plus links that are one standard deviation above the average weight). The visualization shows a clear ring structure mimicking the one observed in networks involving research fields. The connectivity pattern of this network can be interpreted as research fields having a few related activities that share capabilities among them (e.g. capabilities are more re-deployable between molecular biology and biochemistry, than between polymer sciences and experimental psychology). This results in a network structure where each field is connected to a few neighbors.
Finally, we use the same approach to model a “dumbbell” network, which is a network with two well-defined clusters, such as the one observed when connecting skills and occupations [91]. Figures 18 and 19 show an example with 100 economies, 500 activities, and 20 capabilities. We note that obtaining this dumbbell structure requires a good level of mixing between the clusters, which can be achieved by setting the noise levels to be high enough so that some of the between cluster links are comparable in strength to the withing cluster links.
What is exciting about this general idea is that it provides us with an intuitive way to read capability endowments from network structures. For instance, the core periphery-structure of the product space tells us that the capabilities associated with exporting products are correlated among economies, with high complexity economies like that of Singapore, Japan, or the United States, being well endowed with a wide set of capabilities. The ring structure of the research space tells a different story. It is a story of specialization in a world of fine grained capabilities. Similarly, we can use this intuition to think about dumbbell structures, which can be modeled by assuming capability endowments made of slightly overlapping blocks.
@@ 8 Conclusion
Economic complexity has for long attempted to study economic growth and development using methods that are agnostic about the exact nature of factors of production. In this paper, we contribute to this goal by providing an analytical foundation for the economic complexity index ($ECI$) and showing that it can be indeed consider an estimate of the combined presence of undefined or unknown factors of production. For the single-capability model, we could derive the key eigenvector analytically and show that $ECI$ separates economies among those with an above- and below-average probability of having the capability. We then extended this result numerically to a multi-capability setting to show that $ECI$ is a monotonic estimator of an economy’s average capability endowment—even when a substantial share of the capabilities are randomly assigned. In the multi-capability model, $ECI$ is no longer a discrete measure separating low from high capability economies, but a monotonic transformation of the average capability endowment of an economy that recovers the first singular vector of the capability endowment matrix $r_{cb}$. These findings differentiate $ECI$ from measures of diversity, which peak for capability endowments below the maximum (they are non-monotonic functions of $r_{c}$), and thus are non-ideal estimates of the complexity of an economy. These results validate $ECI$ as a measure of composition or complexity, since they show the eigenvector captures information about an economy being endowed with multiple capabilities, regardless of how these capabilities are defined.
@@
Interestingly, our main result does not depend on assuming an stochastic model or a theory based on capabilities, since this idea can be easily generalized to models including factors that are specific to economies and activities. The key condition for the measure of complexity to work is for output to not be perfectly proportional to factor endowments. This condition can be achieved by simply shifting the production function by a constant to make it non-multiplicatively separable. [^24] This mechanism is akin to the idea of symmetry-breaking in physics, since the shift removes symmetries of the function. For example $K_{c}^{\gamma}$ satisfies the scale-invariance symmetry $f(\lambda K)=\lambda^{\gamma}f(K)$, whereas $B+K_{c}^{\gamma}$ lacks this symmetry.
@@
What is also interesting is that the condition needed for our main result to hold comes from calculating the matrix of specialization $R_{cp}$. This is a key difference with previous attempts to connect economic complexity theory and empirics [10, 56] which jumped directly to the binary specialization matrix $M_{cp}$. That assumption results in a monotonic relationship between the number of activities an economy is specialized (its diversity) and its capability endowments [^25] That equation is provided in [56]., which is an uncomfortable result since we have known for a long time that measures of diversity fail to explain future economic growth like measures of complexity do [10]. We now understand that calculating these specialization matrices is a key step, and that skipping this step in theoretical work results in a flawed connection between complexity and capability endowments. This change not only uncovers a tight connection between the economic complexity index and the capability, but explains other findings, like that of Imbs and Warcziag [82], which says that economies diversify only until a certain point.
Our results also open questions about alternative measures of complexity. During the last fifteen years, many alternatives to the economic complexity index have been proposed, such as the Fitness index [19], the Ability index [20], and many others [80, 81, 153, 21, 154, 155, 156, 63]. Since these indexes tend to exhibit strong correlations with $ECI$, our results provide a way to theoretically explore whether they are also monotonic functions of an economy’s capability endowment. In this paper we explore a few alternatives, showing that at least for the cases explored, $ECI$ is a justifiable methodology when output data involves heterogeneous units of observation and noise.
Our work also speaks to the literature attempting to explain the economic complexity index. A key result in this literature is the idea that $ECI$ is a clustering algorithm [58, 62], separating economies into different groups. Our work is consistent with this idea and provides a theoretical interpretation for the clusters, as it shows that what $ECI$ is doing is providing a sigmoid function sorting economies into a high- and low-capability cluster. This sigmoid behavior is a well known feature of the second eigenvector or eigenfunction of diffusion maps, Fokker-Planck operators, and spectral clustering methods (e.g. see [157]). This provides an interesting link to the general idea of diffusion albeit in the context of a model of economic development, opening the door to the notion that these methods could be capturing a generalizable property of economic systems subject to spillovers.
@@
We also embedded this model in a short-run equilibrium framework to estimate wages, consumption, and prices. In this framework, wages increase with capability endowments and prices are higher for products that demand more capabilities. The latter result is highly concave ($\sim 1/(1-q)$), meaning that there is an important premium for producing high complexity products. Interestingly, prices do not strongly affect the specialization condition, meaning that they leave the connection between capability endowments and economic complexity mostly unchanged. [^26] We assume prices are the same across economies.
@@
Together, these findings help resolve a few long-standing tensions in the economic complexity literature. First, and most importantly, the disconnect between its empirical metrics and their theoretical underpinnings. Our findings show that $ECI$ is not an arbitrary or ad-hoc measure, but can be thought of as an estimator of an economy’s capability endowments derived from its pattern of specialization. This is an interesting finding, since it provides a mean to estimate the combined presence of factors or capabilities even when these cannot be identified.
@@
And third, we provide theoretical underpinnings for the structure of the networks of related activities. Our findings show that the structure of these networks is a reflection of how capabilities are distributed across economies and activities.
@@ Acknowledgments
This work owes a very special acknowledgment to Cristian Jara-Figueroa. In 2014, Cristian joined my (César’s) group at the MIT Media Lab. During the first year of his Master’s he worked on the mathematical theory of economic complexity producing an impressive internal manuscript with many results. Those results were never published, but stayed in my group. In 2025, while looking at Cristian’s work, I realized we had made an important and simple mistake at the very beginning, which was to assume that the capability model was a model of $M_{cp}$ instead of a model of the output matrix $Y_{c,p}$. This required reworking everything, and motivated me to estimate the intermediate matrices in the model (such as $R_{cp}$). In my mind this work owes enormously to that effort by Cristian many years ago. We would also like to acknowledge comments by Johanness Wachs, Harasiddh Ganesh, and other members of the Center for Collective Learning. The section on prices and wages was motivated by a inspiring conversation with Jean Tirole.